OpenStudy (clamin):

HELP!!MEDAL! A 20-foot ladder is leaning against a wall. The base of the ladder is 3 feet from the wall. What angle does the ladder make with the ground??

OpenStudy (kittiwitti1):

One question. How far is the base of the ladder from the wall..? The value's missing.

OpenStudy (kittiwitti1):



OpenStudy (clamin):

3

OpenStudy (kittiwitti1):

Okay so put 3 in that diagram. Now from the SOHCAHTOA mnemonic, which trig function do we use on this problem? Sin - opposite and hypotenuse Cos - adjacent and hypotenuse Tan - opposite and adjacent

OpenStudy (clamin):

Cosine

OpenStudy (kittiwitti1):



OpenStudy (kittiwitti1):

Yeah, cosine.

OpenStudy (kittiwitti1):

Do you know where to go from here or are you still lost? :p

OpenStudy (clamin):

I'll try to solve it

OpenStudy (kittiwitti1):

Okay :) I'll be here ^^

OpenStudy (clamin):

Thanks

OpenStudy (kittiwitti1):

? You got it? :O

OpenStudy (clamin):

Is the 15 the angle??

OpenStudy (kittiwitti1):

um... did you arcos the cos fraction?

OpenStudy (kittiwitti1):

arcos is also known as this on the calculator:\[\cos^{-1}\]

OpenStudy (clamin):

Yes and I got 15

OpenStudy (clamin):

Oh wait no my bad

OpenStudy (kittiwitti1):

okay

OpenStudy (kittiwitti1):

Did you get a fraction \[\frac{3}{20}?\]

OpenStudy (clamin):

Yes

OpenStudy (clamin):

I got 81.37

OpenStudy (kittiwitti1):

From this?\[\cos^{-1} \frac{3}{20}\]

OpenStudy (clamin):

Yes

OpenStudy (kittiwitti1):

Okay you have the right answer :) Sorry for the late reply! ^^;

OpenStudy (clamin):

Thanks